OpenStudy (anonymous):

A cannery processed 785 pounds of strawberries in 5.5 hours. The cannery processed 1550 pounds in 10 hours. Write a linear equation to model the relationship between the weight of strawberries S and time T. How many pounds of strawberries can be processed in 8 hours?

OpenStudy (jack1):

all linear equations have the formula: y = mx + c where m = gradient = y2 -y1 / x2 -x1 and c = y intercept

OpenStudy (jack1):

so let y = pounds processed and x = time first solve m = ?

OpenStudy (jack1):

@Bewface u ok with that?

OpenStudy (jack1):

bored now, someone else try?

OpenStudy (anonymous):

Thanks !

OpenStudy (jack1):

hew, so what'd jew get for m?